OpenStudy (anonymous):

t=p+3r/6 solve for r

OpenStudy (anonymous):

it's not that difficult, you're just making r the subject in terms of t and p

OpenStudy (anonymous):

did you get it? so you have: t-p = 3r/6 6(t-p)=3r so r= 6(t-p)/3 which can be simplified to 2(t-p) = r